OpenStudy (anonymous):

what is 12\46 divided by 34\100

Nnesha (nnesha):

\[\large\rm \frac{ \frac{ 12 }{46 } }{ \frac{ 34}{ 100 } }\] like this ??

Nnesha (nnesha):

you can change that to multiplication like this \[\large\rm\frac{ \frac{ a }{ b } }{ \frac{ c }{ d } }= \frac{ a }{ b } \times \frac{ d }{ c }\]

OpenStudy (anonymous):

ok by

Nnesha (nnesha):

got it ?? :)

OpenStudy (anonymous):

yep thx

Nnesha (nnesha):

okay good :) btw \(\huge\color{Green}{{\rm welcome}\rm~to~open~study!!!!}\) \(\Huge \color{blue}{\star^{ \star^{\star:)}}}\Huge \color{red}{\star^{ \star^{\star:)}}}\) \(\Huge \color{orange}{\star^{ \star^{\star:)}}}\Huge \color{purple}{\star^{ \star^{\star:)}}}\)\(\rm\color{green}{o^\wedge\_^\wedge o}\)